Soybean meal is 18 percent protein, cornmeal is 9 percent protein. how many pounds of each should be mixed together in order to get 360 pound mixture that is 10 percent protein?
Answer by mananth(16946)

percent ---------------- quantity
Soyabean 18 ---------------- x pounds
Corn meal 9 ---------------- 360 - x pounds
Mixture 10.00% ---------------- 360

18x+9(360-x)=10 *360

18x+3240-9x=3600
18x-9x=3600-3240
9x= 360
/9
x=40
40 pounds 18.00% Soyabean
320 pounds 9.00% Corn meal
m.ananth@hotmail.ca